Please, help me Simplify: ⁴√64.⁴√32
Answer by rothauserc(4718)

⁴√64.⁴√32 =
16^(1/4) * 4^(1/4) * 16^(1/4) * 2^(1/4) =
note that 16^(1/4) = 2
4 * 8^(1/4) =
4 * 1.681792831 = 6.727171322